OpenStudy (anonymous):

What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0 {(31/8, 25/8)} {(25/8, 31/8)} {(45/2, 31/2)} {(31/2, 45/2)}

OpenStudy (anonymous):

i got {(31/8, 25/8)}

OpenStudy (just_one_last_goodbye):

same @hhelpplzzzz

OpenStudy (trojanpoem):

5x - 3y = 10 x + y = 7 -5x - 5 y = - 35 5 x - 3y = 10 -8y = -25 y = 25/8 x = 7 - 25/8= 31/8 You're correct.

OpenStudy (anonymous):

thank you:)

OpenStudy (mrnood):

well - you can easily check your own answer you should get into habit of doing this just put the values you got into both equations and see if the equations are true